# Subgraph-Aware Training of Text-based Methods for Knowledge Graph Completion

### v3

@@ Appendix E Runtime Analysis
StARWang et al. (2021a) 0.398 0.238 0.487 0.698
@@
StARWang et al. (2021a) 0.288 0.195 0.313 0.480
@@
Table 9: Performance comparison between original StAR and StAR+SaaM where StAR+SaaM stands for the StAR model architecture trained by our training framework SaaM.
Appendix F Effectiveness of SaaM
To demonstrate the generality of SaaM, we apply SaaM to another text-based method StAR (Wang et al., 2021a), which adopted Siamese network architecture. Table 9 shows the results where StAR+SaaM stands for the StAR model architecture trained by our training framework SaaM. Performance improvement are observed in all evaluation metrics, with significant gain in MRR and Hits@1. There results empirically validate that the model-agnostic SaaM framework can be successfully applied to different text-based methods.
Figure 7: The ratio of triples categorized by relation type for the WN18RR, NELL-995, and FB15k-237 datasets, along with the corresponding Hits@1 performance results for each dataset.
Appendix G Performance of Hybrid In-Batch Negatives
@@
Table 10: Performance comparison between Mixed and SaaM, where Mixed replaces half of in-batch triples generated by BRWR with randomly selected triples.
@@
To evaluate the efficacy of constructing a batch exclusively with triples in a subgraph (i.e., SaaM), we compare two sampling methods SaaM and Mixed. Mixed represents a variant that replaces a half of triples in each mini-batch produced by SaaM with randomly selected triples. Table 10 illustrates the performance of SaaM and Mixed. Including random samples deteriorates performances, e.g., a significant drop in Hits@1, which indicates that a mini-batch consisting only of triples within the subgraph is more beneficial.
Triples $\psi_{h}L_{(h,r,t)}$ $L_{(h,r,t)}$
@@
Table 11: Comparison of average loss values on WN18RR for total triples and false positives in the batch.
Approach WN18RR NELL-995 FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Embedding-based approach
SANS Ahrabian et al. (2020) 0.216 0.027 0.322 0.509 0.359 0.309 0.381 0.493 0.298 0.203 0.331 0.486
TransE Bordes et al. (2013) 0.239 0.421 0.450 0.510 0.383 0.321 0.400 0.492 0.280 0.193 0.372 0.439
DistMult Yang et al. (2015) 0.435 0.410 0.450 0.510 0.409 0.343 0.427 0.515 0.280 0.195 0.297 0.441
TuckER Balazevic et al. (2019) 0.466 0.432 0.478 0.518 0.423 0.363 0.455 0.536 0.361 0.265 0.391 0.538
RotatE Sun et al. (2019) 0.471 0.421 0.490 0.568 0.411 0.350 0.439 0.522 0.335 0.243 0.374 0.529
KGTuner Zhang et al. (2022) 0.481 0.438 0.499 0.556 0.428 0.371 0.458 0.544 0.345 0.252 0.381 0.534
CompoundE Ge et al. (2023) 0.492 0.452 0.510 0.570 0.434 0.381 0.466 0.550 0.350 0.262 0.390 0.547
CSProm-KG Chen et al. (2023) 0.569 0.520 0.590 0.675 0.422 0.381 0.474 0.535 0.355 0.261 0.389 0.531
Text-based approach
KG-BERT Yao et al. (2019) 0.216 0.040 0.298 0.516 0.199 0.021 0.267 0.488 0.158 0.019 0.232 0.420
MTL-KGC Kim et al. (2020) 0.331 0.203 0.383 0.597 0.288 0.176 0.229 0.411 0.267 0.172 0.298 0.458
StAR Wang et al. (2021a) 0.398 0.238 0.487 0.698 0.419 0.238 0.395 0.433 0.288 0.195 0.313 0.480
HaSa Zhang et al. (2023a) 0.538 0.444 0.588 0.713 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
KG-S2S Chen et al. (2022) 0.572 0.529 0.595 0.663 0.416 0.311 0.421 0.519 0.337 0.255 0.374 0.496
SimKGC Wang et al. (2022a) 0.671 0.580 0.729 0.811 0.425 0.353 0.447 0.531 0.340 0.252 0.365 0.515
GHN Qiao et al. (2023) 0.678 0.596 0.719 0.821 - - - - 0.339 0.251 0.364 0.518
Ensemble approach
StAR(Ensemble) Wang et al. (2021a) 0.520 0.456 0.509 0.707 0.415 0.311 0.402 0.510 0.332 0.229 0.387 0.526
SATKGC-SPW-DW 0.672 0.605 0.714 0.811 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
SATKGC-DW 0.679 0.609 0.723 0.813 0.429 0.361 0.466 0.551 0.351 0.261 0.372 0.525
SATKGC-SPW 0.681 0.612 0.729 0.820 0.431 0.369 0.472 0.556 0.359 0.271 0.393 0.538
SATKGC 0.689 0.621 0.731 0.823 0.433 0.389 0.476 0.560 0.368 0.276 0.401 0.548
Table 12: KGC results for the WN18RR, NELL-995, and FB15k-237 datasets. “SPW” and “DW” refer to shortest path weight and degree weight respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Qiao et al. (2023).
Appendix H False Positive Analysis
@@
We aim to demonstrate that our contrastive learning with two structure-aware factors selectively penalizes hard negative triples. Table 11 compares $\psi_{h}L_{(h,r,t)}$ and $L_{(h,r,t)}$ on average only for false positives and those for all training triples, where $L_{(h,r,t)}$ in Equation (1) represents loss with only the distance factor $\omega_{h{t_{i}}}$ applied, and $\psi_{h}L_{(h,r,t)}$ in Equation (2) additionally applies the degree factor $\psi_{h}$. The average loss values for the false positives are higher than those for all the triples, which indicates that the structure-aware contrastive learning method severely punishes incorrect predictions for hard negative triples.
Appendix I Correlation of Relation Types and Performances
We investigate the distribution of triples based on their relation types on WN18RR, NELL-995, and FB15k-237. Figure 7 shows that the ratio of triples with the N-N relation type increases in the order of WN18RR, NELL-995, and FB15k-237, while the ratio of triples with the N-1 and 1-N relation types decreases in the same order. In Table 2, both embedding-based and text-based approaches achieve the best results on WN18RR among the three datasets, whereas the performance is worse on FB15k-237. We observe that a high proportion of the N-N relation type and a low proportion of the N-1 and 1-N relation types negatively impact the model performance.
This performance difference between the datasets is larger for a text-based approach than for an embedding-based approach. We speculate that this is because the embedding-based approach randomly initializes the entity and relation embeddings, while the text-based approach uses contextualized text embeddings obtained from PLMs. For the N-N relations where multiple tails can be the correct answer for the same $(h,r)$ pair, the embeddings of these correct tails should be similar. However, PLMs take only text as input, being oblivious of their high similarity. Therefore, these tail embeddings generated by the PLMs might be far apart from each other, so the $(h,r)$ embedding is likely to remain in the middle of these tail embeddings during fine-tuning.
Appendix J Hyperparameter Sensitivity
Figure 8: (a) Impact of varying restart probabilities on the model performance. (b) Impact of varying the number of duplicate entities in a mini-batch on the model performance.
We investigate how restart probability $p_{r}$ in BRWR affects model performance. The hyperparameter $p_{r}$ is associated with the length of the random walk path from the start entity, which in turn influences the occurrence of duplicate entities in a mini-batch. A longer path leads to fewer duplicate entities in the mini-batch. Figure 8(a) illustrates that a lower $p_{r}$ value, encouraging a longer random walk path, leads to higher Hits@1 for WN18RR and FB15k-237. We analyze the impact of duplicate entities in a mini-batch on the model performance. In Figure 8(b), more duplicate entities resulting from higher $p_{r}$ negatively impact on the performance, which highlights the importance of reducing the duplicates in a mini-batch to avoid the performance degradation.
Appendix K Entire Main Results

### v5

@@ Appendix G Performance of the SaaM Scheme Across Models
RotatE(Sun et al., 2019) 0.471 0.421 0.490 0.568
RotatE+SaaM 0.479 0.433 0.505 0.580
StAR(Wang et al., 2021b) 0.398 0.238 0.487 0.698
@@
RotatE(Sun et al., 2019) 0.335 0.243 0.374 0.529
RotatE+SaaM 0.343 0.255 0.389 0.534
StAR(Wang et al., 2021b) 0.288 0.195 0.313 0.480
@@
Table 8. Performance comparison between original StAR and StAR+SaaM where StAR+SaaM stands for the StAR model architecture trained by our training framework SaaM.
@@
Table 9. Performance comparison between Mixed and SaaM, where Mixed replaces half of in-batch triples generated by SaaM with randomly selected triples.
Figure 10. The proportion of triples categorized by relation types for the Wikidata5M-Trans, WN18RR, and FB15k-237 datasets, along with their corresponding Hits@1 performance results.
Preprocessing Time
Algorithm WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
BRWR 8m 12m 12h 12h
Shortest Path + Degree 7m 10m 7h 7h
Training Time Per Epoch
Model WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
SATKGC 4m 6m 10h 10h
SimKGC (Wang et al., 2022) 3m 5m 9h 9h
StAR (Wang et al., 2021b) 1h 1h 30m - -
Table 10. The elapsed time required for sampling and the time per epoch during training.
Distance-Similarity Analysis
Distance 1 2 3 4 5 6 7 8
Avg. Cosine Similarity 0.818 0.762 0.611 0.323 0.152 -0.04 -0.219 -0.378
Table 11. Average cosine similarity between two entities at their varying distances in FB15k-237. The decreasing similarity scores with increasing distances demonstrate that our model effectively captures the structural properties of the knowledge graph.
Appendix H Performance of Hybrid In-Batch Negatives
@@
To evaluate the efficacy of constructing a batch exclusively from triples in a subgraph (i.e., SaaM), we compare two sampling methods: SaaM and Mixed. Mixed is a variant that replaces half of the triples in each mini-batch generated by SaaM with randomly selected triples. Table 9 illustrates the performance of SaaM and Mixed. Including randomly selected triples degrades performance, as evidenced by a significant drop in Hits@1, indicating that a mini-batch composed solely of triples within the subgraph is more beneficial.
Appendix I Correlation of Relation Types and Performances
We investigate the distribution of triples based on their relation types on Wikidata5M-Trans, WN18RR, and FB15k-237. Figure 10 shows that the proportion of triples with the N-N relation type increases in the order of Wikidata5M-Trans, WN18RR, and FB15k-237, while the proportion of triples with the N-1 relation type decreases in the same order. In Table 2 and Table 3, We observe that text-based methods outperform embedding-based methods in Wikidata and WN18RR, which have a higher proportion of N-1 relation type and lower proportion of N-N relation type. In contrast, in FB15k-237, which has a higher proportion of N-N relation type and a lower proportion of N-1 relation type, embedding-based methods generally achieve better performance than text-based methods. This performance difference between the datasets is larger for a text-based approach than for an embedding-based approach. We speculate that this is because the embedding-based approach randomly initializes the entity and relation embeddings, while the text-based approach uses contextualized text embeddings obtained from PLMs. For the N-N relations where multiple tails can be the correct answer for the same $(h,r)$ pair, the embeddings of these correct tails should be similar. However, PLMs take only text as input, being oblivious of their high similarity. Therefore, these tail embeddings generated by the PLMs might be far apart from each other, so the $(h,r)$ embedding is likely to remain in the middle of these tail embeddings during fine-tuning.
Appendix J Runtime Analysis
Running SATKGC incurs a marginal computational overhead, because (i) sampling subgraphs and (ii) computing distances and degrees are performed in advance before training. As shown in Table 10, the computational cost for (i) and (ii) is acceptable, and depends on the mini-batch size, which can be adjusted. Moreover, the time complexity for (ii) is acceptable. For each mini-batch $B$ of triples, we run Dijkstra’s single source shortest path algorithm, and thus the runtime to compute the distance is O($|V|log|V|+|B|log|V|$), where $V$ is a set of entities in $B$ [^15] As described in Section 4.3, we do not calculate all-pair shortest paths for every pair of vertices in $V$. To reduce the computational overhead, we compute the approximate distance by using the shortest path between the head of the center triple in a subgraph and every tail in that subgraph, which have been already obtained fro Dijkstra’s algorithm. Table 10 shows the training time per epoch for SATKGC, SimKGC (Wang et al., 2022), and StAR (Wang et al., 2021b). SATKGC remains competitive, though it takes slightly more time than SimKGC due to the computational overhead for (a) computing its loss using the shortest path length and the degree, (b) counting the occurrences of visited triples in SaaM, and (c) fetching subgraphs in SaaM. [^16] StAR runs out of memory on the Wikidata5M datasets.
Appendix K Model Representation Analysis
We analyze how effectively SATKGC captures both textual and structural properties of the knowledge graph. To evaluate its textual representation capabilities, we examine a variant of SATKGC without Proximity-aware Contrastive Learning and Frequency-aware Mini-batch Training. This model achieves Hit@1 scores of 0.608 and 0.201 on FB15k237 and WN18RR, respectively, demonstrating strong textual representation, owing to the expressive power of the underlying PLM.
To assess whether our method preserves the structural properties of the knowledge graph, we analyze the relation of the distance between two entities in the knowledge graph and their representational similarity. For this, we compute the average cosine similarity between every two entities’ representations by varying their distance (ranging from 1 to 8) in FB15k-237. The results presented in Table 11 reveal a clear negative correlation between the distance and the similarity, i.e., the similarity score consistently decreases as the distance increases, ranging from 0.818 for directly connected entities to -0.378 for entities reached from each other by eight hops. This clear decline indicates that SATKGC effectively encodes structural properties by assigning lower similarity scores to more distant entities.
Approach WN18RR FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Embedding-based approach
SANS (Ahrabian et al., 2020) 0.216 0.027 0.322 0.509 0.298 0.203 0.331 0.486
TransE (Bordes et al., 2013) 0.239 0.421 0.450 0.510 0.280 0.193 0.372 0.439
DistMult (Yang et al., 2015) 0.435 0.410 0.450 0.510 0.280 0.195 0.297 0.441
TuckER (Balazevic et al., 2019) 0.466 0.432 0.478 0.518 0.361 0.265 0.391 0.538
RotatE (Sun et al., 2019) 0.471 0.421 0.490 0.568 0.335 0.243 0.374 0.529
KGTuner (Zhang et al., 2022) 0.481 0.438 0.499 0.556 0.345 0.252 0.381 0.534
KG-Mixup (Shomer et al., 2023) 0.488 0.443 0.505 0.541 0.359 0.265 0.395 0.547
[…]
CompoundE (Ge et al., 2023) 0.492 0.452 0.510 0.570 0.350 0.262 0.390 0.547
KRACL (Tan et al., 2023) 0.529 0.480 0.539 0.621 0.360 0.261 0.393 0.548
CSProm-KG (Chen et al., 2023) 0.569 0.520 0.590 0.675 0.355 0.261 0.389 0.531
Text-based approach
KG-BERT (Yao et al., 2019) 0.216 0.040 0.298 0.516 0.158 0.019 0.232 0.420
MTL-KGC (Kim et al., 2020) 0.331 0.203 0.383 0.597 0.267 0.172 0.298 0.458
StAR (Wang et al., 2021b) 0.398 0.238 0.487 0.698 0.288 0.195 0.313 0.480
HaSa (Zhang et al., 2023c) 0.535 0.446 0.587 0.711 0.301 0.218 0.325 0.482
KG-S2S (Chen et al., 2022) 0.572 0.529 0.595 0.663 0.337 0.255 0.374 0.496
SimKGC (Wang et al., 2022) 0.671 0.580 0.729 0.811 0.340 0.252 0.365 0.515
[…]
Ensemble approach
StAR(Self-Adp) (Wang et al., 2021b) 0.520 0.456 0.509 0.707 0.332 0.229 0.387 0.526
SATKGC (w/o SaaM) 0.673 0.595 0.728 0.813 0.349 0.256 0.367 0.520
SATKGC (w/o PCL, FMT) 0.676 0.608 0.722 0.820 0.355 0.261 0.386 0.537
SATKGC (w/o FMT) 0.680 0.611 0.729 0.823 0.351 0.265 0.389 0.541
SATKGC (w/o PCL) 0.686 0.623 0.740 0.827 0.366 0.272 0.396 0.545
SATKGC 0.694 0.626 0.743 0.833 0.370 0.279 0.403 0.550
Table 12. KGC results for the WN18RR, and FB15k-237 datasets. “PCL” and “FMT” refer to Proximity-aware Contrastive Learning and Frequency-aware Mini-batch Training respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Liu et al. (2024) $\ddagger$: numbers are from Qiao et al. (2023).
Triples $\psi_{\mathcal{G}}(t)\mathcal{L}_{(h,r,t)}$ $\mathcal{L}_{(h,r,t)}$
@@ Appendix I Correlation of Relation Types and Performances
Table 13. Comparison of average loss values on WN18RR for total triples and false positives in the batch.
Appendix L False Positive Analysis
@@
We aim to demonstrate that our contrastive learning with two structure-aware factors selectively penalizes hard negative triples. Table 13 compares $\psi_{\mathcal{G}}(t)\mathcal{L}_{(h,r,t)}$ and $\mathcal{L}_{(h,r,t)}$ on average, specifically for false positives of the existing state-of-the-art model (Wang et al., 2022) and for all training triples. In this comparison, $\mathcal{L}_{(h,r,t)}$ in Equation (3) represents the loss with only a structural hardness factor $\omega_{h{t_{i}}}$ applied, while $\psi_{\mathcal{G}}(t)\mathcal{L}_{(h,r,t)}$ in Equation (5) additionally incorporates a reweighting factor $\psi_{\mathcal{G}}(t)$. The average loss values for the false positives are higher than those for all the triples, which indicates that the structure-aware contrastive learning method severely punishes incorrect predictions for hard negative triples.
Wikidata5M-Trans
Model MRR Hits@1 Hits@3 Hits@10
ReSKGC 0.396 0.373 0.413 0.437
SATKGC 0.408 0.366 0.425 0.479
Table 14. Performance comparison between SATKGC and ReSKGC. ReSKGC is a retrieval-augmented generation(RAG)-based model that performs search on the knowledge graph rather than learning the knowledge graph itself.
Appendix M Peformance Comparison with RAG-based Model
We additionally conduct performance comparison experiments with the RAG-based model ResKGC (Yu and Yang, 2023). Since ReSKGC does not have publicly available source code, we use the performance results reported directly in the paper. Furthermore, as there are no reported results for FB15k-237 and WN18RR, we only present the results on Wikidata5M-Trans. As shown in Table 14, SATKGC demonstrates superior performance across all metrics except for Hits@1. Unlike SATKGC, ReSKGC is a decoder-based generation model. When given the head and relation IDs, it converts them into text, retrieves the most semantically relevant triples from the KG using a retrieval component such as BM25, and generates the correct tail based on that information. Therefore, unlike SATKGC, ReSKGC does not learn the KG directly and only uses it for retrieval, without utilizing any structural information of the KG. This explains why SATKGC achieves overwhelmingly better performance, despite both models having the same number of parameters used during training.
Appendix N Entire Main Results